# Quantised helicity in optical media

### v1

@@ Abstract
Discussion. In this Letter, we set out to find an expression for the optical helicity which is valid within any optical medium, and to resolve the mutual incompatibility between the duality transform (Quantised helicity in optical media), and the constitutive equations. We approached this by using the duality transform (Quantised helicity in optical media) as the starting point from which to define the optical helicity, which provides the appropriate rotation between $\mathbf{E}$ and $\mathbf{H}$, and also $\mathbf{D}$ and $\mathbf{B}$, without coming into conflict with constitutive relations or presupposing any particular properties of the medium. This works because, in our approach, the constitutive equations connecting the fields emerge as a consequence of the Lagrangian and the associated equations of motion, rather than being enforced from the outset. We therefore require a description of the material’s internal degrees of freedom in the Lagrangian, and this explicit description of the material causes a medium contribution to the helicity to naturally emerge.
The extension of the Lagrangian (18) to allow for multiple resonances in $\varepsilon$ and $\mu$ may be achieved by straightforward modification of the harmonic oscillator terms. Lossy media can be treated by coupling the $\mathbf{P}$ and $\mathbf{M}$ fields to a continuous reservoir of harmonic oscillators [26, 27, 28, 29, 30]. Finally, chiral and magnetoelectric responses can also be added to these models by introducing additional terms into $\mathcal{L}$ which couple the $\mathbf{P}$ and $\mathbf{M}$ degrees of freedom [31]. As we have argued, any such modifications will not affect the basic results (9) and (14).
@@
When comparing our results to those obtained elsewhere in the literature, it is important to appreciate that (Quantised helicity in optical media) and (Quantised helicity in optical media) are, in general, distinct symmetry transformations, and so one would not expect the conserved quantities associated with these to coincide exactly. Even in the case where $\varepsilon=\mu$, in which (Quantised helicity in optical media) reduces to (Quantised helicity in optical media), we note that the helicity arising from Eq. (Quantised helicity in optical media) contains a matter contribution absent from the helicity derived from Eq. (Quantised helicity in optical media). This situation is analogous to that discussed in Ref. [23] which points out that the total electromagnetic energy used in phenomenological quantisation schemes differs from the total energy of the full system by a matter component.
@@
A point worth mentioning is that the definition presented here relies on the ability to use a model for the medium that can be made symmetrical under the duality transform (Quantised helicity in optical media), for at least some choice of model parameters. This requires that the polarisation and magnetisation can be written as functions of $\mathbf{D}$ and $\mathbf{B}$, or $\mathbf{E}$ and $\mathbf{H}$, but not some combination of the two. For example, a medium where $\mathbf{P}=\mathbf{P}(\mathbf{E},\mathbf{H})$ and $\mathbf{M}=\mathbf{M}(\mathbf{H},\mathbf{E})$ would fulfil this requirement, but one where $\mathbf{P}=\mathbf{P}(\mathbf{E},\mathbf{B})$ and $\mathbf{M}=\mathbf{M}(\mathbf{B},\mathbf{E})$ would not. It is possible to reformulate the description of such materials in a way that can be made symmetric under the duality transform (Quantised helicity in optical media); we leave this to future work.
As a final point, we may note that, unlike the helicity density $h$, the helicity flux density, $\mathbf{v}$, derived from our procedures contains no additional medium component. This makes physical sense, as the medium constituents cannot themselves flow, leaving the task of carrying helicity from place to place entirely to the electromagnetic fields.
@@ Acknowledgements.
Acknowledgements. This research was supported by the Royal Commission for the Exhibition of 1851, the Royal Society (grant numbers RSRP/RE/220013 and RP/150122), the EPSRC (grant number EP/V048449/1) and the Leverhulme Trust.
Supplemental Information for
Quantised helicity in optical media
I Noether’s Theorem
Here we provide the details of the application of Noether’s theorem [32] to derive an expression for the optical helicity from the Lagrangian (18). Under the infinitesimal version of duality transformation (Quantised helicity in optical media),
@@ I Noether’s Theorem
$\mathbf{P}^{\prime}=\mathbf{P}+\delta\theta\mathbf{M},$ $\mathbf{M}^{\prime}=\mathbf{M}-\delta\theta\mathbf{P},$ (SI 1)
the Lagrangian becomes $\mathcal{L}^{\prime}=\mathcal{L}(\mathbf{A^{\prime}},\mathbf{C^{\prime}},\mathbf{P^{\prime}},\mathbf{M^{\prime}})$. We first consider the change in the Lagrangian, $\delta\mathcal{L}\equiv\mathcal{L}^{\prime}-\mathcal{L}$. Substituting the duality transformation into the Lagrangian (18), we find that
$\delta\mathcal{L}$ $=\delta\theta\left(\mathbf{E}\cdot\mathbf{B}+\mathbf{H}\cdot\mathbf{D}\right)$ (SI 2)
@@
[…]
where we have used that $\boldsymbol{\Pi}^{P}=\partial_{t}\mathbf{P}/\alpha^{2}$ and $\boldsymbol{\Pi}^{M}=\partial_{t}\mathbf{M}/\alpha^{2}$ in the chosen gauge and coupling. Equating the two expressions for $\delta\mathcal{L}$ from Eq. (SI 2) and Eq. (SI 3) leads to the local conservation law
$+\boldsymbol{\nabla}\cdot\left[\frac{1}{2}\left(\mathbf{E}\times\mathbf{A}+\mathbf{H}\times\mathbf{C}\right)\right]=0,$ (SI 4)
II Hamiltonian in field variables
If we consider a dual-symmetric medium (that is, we set $\alpha=\beta$ and $\omega_{E}=\omega_{M}=\omega_{0}$), then the Hamiltonian obtained from the Lagrangian in Eq. (18) takes the form
$\hat{H}$ $=\int d^{3}r\;\frac{1}{2}\left[\left(\boldsymbol{\widehat{\Pi}}^{A}\right)^{2}+\mathbf{\widehat{B}}^{2}\right]$
$\hskip 42.67912pt+\frac{1}{2\alpha^{2}}\left[\alpha^{4}\left(\boldsymbol{\widehat{\Pi}}^{P}\right)^{2}+\left(\omega_{0}^{2}+\alpha^{2}\right)\mathbf{\widehat{P}}^{2}\right]$
$\hskip 42.67912pt+\frac{1}{2\alpha^{2}}\left[\alpha^{4}\left(\boldsymbol{\widehat{\Pi}}^{M}\right)^{2}+\left(\omega_{0}^{2}+\alpha^{2}\right)\mathbf{\widehat{M}}^{2}\right]$
which is dual symmetric with respect to the transform in $(\mathbf{\widehat{D}},\mathbf{\widehat{B}})$, as $\boldsymbol{\widehat{\Pi}}^{A}=-\mathbf{\widehat{D}}$. If we expand the fields as
$\mathbf{\widehat{P}}$ $=\sum_{\lambda=\pm}\int\frac{d^{3}k}{(2\pi)^{3/2}}\;\frac{\alpha}{\sqrt{2\widetilde{\omega}_{0}}}\bigg{[}\mathbf{e}_{\lambda}(\mathbf{k})\hat{b}_{\lambda}(\mathbf{k})e^{i\mathbf{k}\cdot\mathbf{x}}+\mathbf{e}_{-\lambda}(\mathbf{k})\hat{b}^{\dagger}_{\lambda}(\mathbf{k})e^{-i\mathbf{k}\cdot\mathbf{x}}\bigg{]},$ (SI 5)
$\mathbf{\widehat{M}}$ $=\sum_{\lambda=\pm}\int\frac{d^{3}k}{(2\pi)^{3/2}}\;\frac{\alpha}{\sqrt{2\widetilde{\omega}_{0}}}\bigg{[}\mathbf{e}_{\lambda}(\mathbf{k})\hat{c}_{\lambda}(\mathbf{k})e^{i\mathbf{k}\cdot\mathbf{x}}+\mathbf{e}_{-\lambda}(\mathbf{k})\hat{c}^{\dagger}_{\lambda}(\mathbf{k})e^{-i\mathbf{k}\cdot\mathbf{x}}\bigg{]},$
then we arrive at the Hamiltonian seen in the main text (22).
III Details of the diagonalisation
Here we present the details of the diagonalisation of (22), leading to expressions for the Hopfield coefficients appearing in (Quantised helicity in optical media). We begin with some preparatory results, which will be useful in expanding the helicity operator to obtain Eq. (Quantised helicity in optical media) . First, we can invert Eq. (Quantised helicity in optical media) by noting that, as the polariton operators form a complete basis set, we must have that $\hat{a}_{\lambda}(\mathbf{k})=\sum_{i}f_{i}\hat{p}_{i}(\mathbf{k},\lambda)+g_{i}\hat{p}^{\dagger}_{i}(-\mathbf{k},\lambda)$ for some constants $f_{i}$ and $g_{i}$ (and similarly for the other operators). Computing the commutators $\left[\hat{a}_{\lambda}(\mathbf{k}^{\prime}),\hat{p}_{i}(\mathbf{k},\lambda)\right]$ and $\left[\hat{a}_{\lambda}(\mathbf{k}^{\prime}),\hat{p}^{\dagger}_{i}(-\mathbf{k},\lambda)\right]$ allows us to identify the constants,
@@ III Details of the diagonalisation
$\hat{b}_{\lambda}(\mathbf{k})$ $=\sum_{i}\beta_{i}^{*}\hat{p}_{i}(\mathbf{k},\lambda)-\widetilde{\beta}_{i}\hat{p}^{\dagger}_{i}(-\mathbf{k},\lambda),$ (SI 6)
@@
Furthermore, inserting the expressions in Eqns. (SI 6) into the commutators $\left[\hat{a}_{\lambda}(\mathbf{k}),\hat{a}^{\dagger}_{\lambda^{\prime}}(\mathbf{k}^{\prime})\right]=\delta_{\lambda\lambda^{\prime}}\delta(\mathbf{k}-\mathbf{k}^{\prime})$, etc., produces the relations
@@
$\sum_{i}|\beta_{i}|^{2}-|\widetilde{\beta}_{i}|^{2}$ $=1,$ (SI 7)
@@
$(\omega_{i}-k)\alpha_{i}$ $=\Lambda\left[-i(\beta_{i}-\widetilde{\beta}_{i})-\lambda(\gamma_{i}-\widetilde{\gamma}_{i})\right],$
$(\omega_{i}+k)\widetilde{\alpha}_{i}$ $=\Lambda\left[i(\beta_{i}-\widetilde{\beta}_{i})-\lambda(\gamma_{i}-\widetilde{\gamma}_{i})\right],$
$\beta_{i}$ $=i\Lambda\left(\frac{\alpha_{i}+\widetilde{\alpha}_{i}}{\omega_{i}-\widetilde{\omega}_{0}}\right),$
$\widetilde{\beta}_{i}$ $=i\Lambda\left(\frac{\alpha_{i}+\widetilde{\alpha}_{i}}{\omega_{i}+\widetilde{\omega}_{0}}\right),$ (SI 8)
$\gamma_{i}$ $=-\lambda\Lambda\left(\frac{\alpha_{i}-\widetilde{\alpha}_{i}}{\omega_{i}-\widetilde{\omega}_{0}}\right),$
$\widetilde{\gamma}_{i}$ $=-\lambda\Lambda\left(\frac{\alpha_{i}-\widetilde{\alpha}_{i}}{\omega_{i}+\widetilde{\omega}_{0}}\right).$
@@
$|\alpha_{i}|^{2}-|\widetilde{\alpha}_{i}|^{2}+|\beta_{i}|^{2}-|\widetilde{\beta}_{i}|^{2}+|\gamma_{i}|^{2}-|\widetilde{\gamma}_{i}|^{2}=1.$ (SI 9)
@@
$n(\omega_{i})=\frac{k}{\omega_{i}}=\left(1-\frac{\alpha^{2}}{\omega_{i}^{2}-\omega_{0}^{2}}\right),$ (SI 10)
@@
$n(\omega_{i})=\frac{k}{\omega_{i}}=-\left(1-\frac{\alpha^{2}}{\omega_{i}^{2}-\omega_{0}^{2}}\right),$ (SI 11)
@@
where we have chosen the modes $\omega_{i}=\{\omega_{u},\omega_{p},\omega_{d}\}$ in descending order of oscillation frequency. Manipulating the algebratic relations in Eqns. (III) yields
@@
$\left|\alpha_{u/d}\right|^{2}$ $=\left(1+2\Lambda^{2}\left[\frac{1}{(\omega_{u/d}-\widetilde{\omega}_{0})^{2}}-\frac{1}{(\omega_{u/d}+\widetilde{\omega}_{0})^{2}}\right]\right)^{-1}$ (SI 12)
$\left|\widetilde{\alpha}_{p}\right|^{2}$ $=\left(-1+2\Lambda^{2}\left[\frac{1}{(\omega_{p}-\widetilde{\omega}_{0})^{2}}-\frac{1}{(\omega_{p}+\widetilde{\omega}_{0})^{2}}\right]\right)^{-1}$ (SI 13)
using Eqns. (III) and the normalisation condition Eq. (SI 9). Importantly, we can note that this satisfies
@@
which must be the case in order to preserve the canonical commutation relations. The phases of $\alpha_{u/d}$ and $\widetilde{\alpha}_{p}$ can be chosen freely, and the remaining Hopfield coefficients follow from Eqns. (III).
IV Diagonal form of the helicity
In this section, we provide details of how to obtain Eq. (Quantised helicity in optical media) in the main text, using the polariton coefficients obtained in the previous section. Let us express the total helicity, Eq. (9), in terms of the photon and matter annihilation and creation operators, yielding:
$\widehat{\mathcal{H}}$ $=\int d^{3}r\;\left[\frac{1}{2}\left(\mathbf{\widehat{A}}\cdot\mathbf{\widehat{B}}-\mathbf{\widehat{C}}\cdot\mathbf{\widehat{D}}\right)+\boldsymbol{\widehat{\Pi}}^{P}\cdot\mathbf{\widehat{M}}-\boldsymbol{\widehat{\Pi}}^{M}\cdot\mathbf{\widehat{P}}\right]$

### v2

@@ VI Discussion and conclusion
In this article, we set out to find an expression for the optical helicity which is valid within any optical medium, and to resolve the mutual incompatibility between the duality transform (I), and the constitutive equations. We approached this by using the duality transform (II) as the starting point from which to define the optical helicity, which provides the appropriate rotation between $\mathbf{E}$ and $\mathbf{H}$, and also $\mathbf{D}$ and $\mathbf{B}$, without coming into conflict with constitutive relations or presupposing any particular properties of the medium. This works because, in our approach, the constitutive equations connecting the fields emerge as a consequence of the Lagrangian and the associated equations of motion, rather than being enforced from the outset. We therefore require a description of the material’s internal degrees of freedom in the Lagrangian, and this explicit description of the material causes a medium contribution to the helicity to naturally emerge.
We find that the helicity defined in this way retains many properties that we are familiar with, but differences arise as this helicity also contains the twist associated with the matter. As polaritons are the natural quasi-particle excitations in optical media, we have used them to frame our discussions. In all cases, when the polariton is photon-like, it retains the interpretation of carrying $\pm$ the energy density over frequency. However, it is generally the total energy that is important, rather than only the electromagnetic energy. We find that, in a dual-symmetric medium where $\varepsilon=\mu$, the helicity per polariton is always $\pm$ the total energy density over polariton frequency. However, in more general optical media, the helicity is not conserved, leading to a $k$-varying helicity for single polariton-branch excitations. Broadly speaking, the $k$-dependence can be seen as a consequence of the fact that, unlike in free space or in a dual-symmetric medium, the field vectors for circularly polarised modes are not generally eigenstates of the duality transformation; this point of view is outlined mathematically in Appendix F. The precise details of the $k$-dependence depend on the polariton character, and the details of the polarisation and magnetisation response. Interestingly, this non-conservation also leads to the possibility of helicity oscillations, akin to neutrino oscillations.
@@
When comparing our results to those obtained elsewhere in the literature, it is important to appreciate that (II) and (II) are, in general, distinct symmetry transformations, and so one would not expect the conserved quantities associated with these to coincide exactly. Even in the case where $\varepsilon=\mu$, in which (II) reduces to (II), we note that the helicity arising from Eq. (II) contains a matter contribution absent from the helicity derived from Eq. (II). This situation is analogous to that discussed in Ref. [26] which points out that the total electromagnetic energy used in phenomenological quantisation schemes differs from the total energy of the full system by a matter component. Both are, of course, valid notions of helicity. We argue, however, that the approach taken here is superior as it explicitly takes the contribution from the matter into account. A somewhat different way of defining a “matter” contribution to the helicity has been presented by Fernandez-Corbaton [33], where the static, zero-frequency contribution to the helicity in the presence of magnetic dipole sources is interpreted as helicity stored by the matter.
The electromagnetic helicity and optical chirality [34] are closely related quantities, and it has been proposed to generalise the optical chirality to dispersive, lossy media by taking $\frac{1}{2}\Big(\mathbf{E}\times(\boldsymbol{\nabla}\times\mathbf{H})+\mathbf{H}\times(\boldsymbol{\nabla}\times\mathbf{E})\Big)$ as the chirality flux in a medium [35]. For monochromatic fields in free space, the chirality density is equal to $\omega^{2}$ times the helicity density, and likewise for the fluxes, and it might be supposed that a similar relationship should hold for monochromatic fields in a lossless medium. A similar relationship between the fluxes does indeed hold, but only for a different possible definition of the chirality flux: our helicity flux, $\mathbf{v}$, is given by (17), and so a chirality flux that is equal to $(\omega^{2}/v_{p}^{2})\mathbf{v}$ for monochromatic fields would be $\frac{1}{2}\Big(\mathbf{E}\times(\boldsymbol{\nabla}\times\mathbf{B})+\mathbf{H}\times(\boldsymbol{\nabla}\times\mathbf{D})\Big)$.
The extension of the Lagrangian (21) to allow for multiple resonances in $\varepsilon$ and $\mu$ may be achieved by straightforward modification of the harmonic oscillator terms. Lossy media can be treated by coupling the $\mathbf{P}$ and $\mathbf{M}$ fields to a continuous reservoir of harmonic oscillators [36, 37, 38, 39, 40, 41]. Finally, chiral and magnetoelectric responses can also be added to these models by introducing additional terms into $\mathcal{L}$ which couple the $\mathbf{P}$ and $\mathbf{M}$ degrees of freedom [42]. Such changes will generally alter the expressions for the conjugate momenta $\mathbf{\Pi}^{P}$ and $\mathbf{\Pi}^{M}$ (22-23), and their interpretation. However, any such modifications will not affect the basic results (9) and (17).
@@
A point worth mentioning is that the definition presented here relies on the ability to use a model for the medium that can be made symmetrical under the duality transform (II), for at least some choice of model parameters. This requires that the polarisation and magnetisation can be written as functions of $\mathbf{D}$ and $\mathbf{B}$, or $\mathbf{E}$ and $\mathbf{H}$, but not some combination of the two. For example, a medium where $\mathbf{P}=\mathbf{P}(\mathbf{E},\mathbf{H})$ and $\mathbf{M}=\mathbf{M}(\mathbf{H},\mathbf{E})$ would fulfil this requirement, but one where $\mathbf{P}=\mathbf{P}(\mathbf{E},\mathbf{B})$ and $\mathbf{M}=\mathbf{M}(\mathbf{B},\mathbf{E})$ would not. It is possible to reformulate the description of such materials in a way that can be made symmetric under the duality transform (II); we leave this to future work.
@@ Acknowledgements.
This research was supported by the Royal Commission for the Exhibition of 1851, the Royal Society (grant numbers RSRP/RE/220013, RP/150122 and RSRP/R/210005), the EPSRC (grant number EP/V048449/1) and the Leverhulme Trust.
Appendices
Appendix A Noether’s Theorem
Here we provide the details of the application of Noether’s theorem [43] to derive an expression for the optical helicity from the Lagrangian (21). Under the infinitesimal version of duality transformation (II),
@@ Appendix A Noether’s Theorem
$\mathbf{P}^{\prime}=\mathbf{P}+\delta\theta\mathbf{M},$ $\mathbf{M}^{\prime}=\mathbf{M}-\delta\theta\mathbf{P},$ (39)
the Lagrangian becomes $\mathcal{L}^{\prime}=\mathcal{L}(\mathbf{A^{\prime}},\mathbf{C^{\prime}},\mathbf{P^{\prime}},\mathbf{M^{\prime}})$. We first consider the change in the Lagrangian, $\delta\mathcal{L}\equiv\mathcal{L}^{\prime}$-$\mathcal{L}$. Substituting the duality transformation into the Lagrangian (21), and setting $\alpha=\beta$ and $\omega_{E}=\omega_{M}$, we find that
$\delta\mathcal{L}$ $=\delta\theta\left(\mathbf{E}\cdot\mathbf{B}+\mathbf{H}\cdot\mathbf{D}\right)$ (40)
@@
We then calculate $\delta S=\int_{t_{A}}^{t_{B}}dt\;\delta\mathcal{L}$ through variation of the fields. The result is given by
$\delta S$ $=\delta\theta\int d^{3}r\;\int_{t_{A}}^{t_{B}}dt\left(\frac{\partial}{\partial t}\left[-\mathbf{C}\cdot\mathbf{D}+\boldsymbol{\Pi}^{P}\cdot\mathbf{M}-\boldsymbol{\Pi}^{M}\cdot\mathbf{P}\right]-\boldsymbol{\nabla}\cdot\left[\mathbf{H}\times\mathbf{C}\right]\right),$ (41)
where we have used that $\boldsymbol{\Pi}^{P}=\partial_{t}\mathbf{P}/\alpha^{2}$ and $\boldsymbol{\Pi}^{M}=\partial_{t}\mathbf{M}/\alpha^{2}$ in the chosen gauge and coupling. Equating the two expressions for $\delta\mathcal{L}$ from Eq. (40) and Eq. (41) leads to the local conservation law
$+\boldsymbol{\nabla}\cdot\left[\frac{1}{2}\left(\mathbf{E}\times\mathbf{A}+\mathbf{H}\times\mathbf{C}\right)\right]=0,$ (42)
Appendix B Hamiltonian in field variables
In terms of $\mathbf{A}$, $\mathbf{P}$ and $\mathbf{M}$, and their conjugate variables, the Hamiltonian obtained from the Lagrangian in Eq. (21) takes the form
$\hat{H}$ $=\int d^{3}r\;\frac{1}{2}\left[\left(\boldsymbol{\widehat{\Pi}}^{A}\right)^{2}+\mathbf{\widehat{B}}^{2}\right]+\frac{1}{2\alpha^{2}}\left[\alpha^{4}\left(\boldsymbol{\widehat{\Pi}}^{P}\right)^{2}+\widetilde{\omega}_{E}^{2}\mathbf{\widehat{P}}^{2}\right]+\frac{1}{2\beta^{2}}\left[\beta^{4}\left(\boldsymbol{\widehat{\Pi}}^{M}\right)^{2}+\widetilde{\omega}_{M}^{2}\mathbf{\widehat{M}}^{2}\right]$
If we expand the fields as
$\mathbf{\widehat{P}}$ $=\sum_{\lambda=\pm}\int\frac{d^{3}k}{(2\pi)^{3/2}}\;\frac{\alpha}{\sqrt{2\widetilde{\omega}_{E}}}\bigg[\mathbf{e}_{\lambda}(\mathbf{k})\hat{b}_{\lambda}(\mathbf{k})e^{i\mathbf{k}\cdot\mathbf{x}}+\mathbf{e}_{-\lambda}(\mathbf{k})\hat{b}^{\dagger}_{\lambda}(\mathbf{k})e^{-i\mathbf{k}\cdot\mathbf{x}}\bigg],$ (43)
$\mathbf{\widehat{M}}$ $=\sum_{\lambda=\pm}\int\frac{d^{3}k}{(2\pi)^{3/2}}\;\frac{\beta}{\sqrt{2\widetilde{\omega}_{M}}}\bigg[\mathbf{e}_{\lambda}(\mathbf{k})\hat{c}_{\lambda}(\mathbf{k})e^{i\mathbf{k}\cdot\mathbf{x}}+\mathbf{e}_{-\lambda}(\mathbf{k})\hat{c}^{\dagger}_{\lambda}(\mathbf{k})e^{-i\mathbf{k}\cdot\mathbf{x}}\bigg],$
@@ Appendix B Hamiltonian in field variables
Appendix C Details of the diagonalisation
Here we present the details of the diagonalisation of (26), leading to expressions for the Hopfield coefficients appearing in (V.1). We begin with some preparatory results, which will be useful in expanding the helicity operator to obtain Eq. (29) . First, we can invert Eq. (V.1) by noting that, as the polariton operators form a complete basis set, we must have that $\hat{a}_{\lambda}(\mathbf{k})=\sum_{i}f_{i}\hat{p}_{i}(\mathbf{k},\lambda)+g_{i}\hat{p}^{\dagger}_{i}(-\mathbf{k},\lambda)$ for some constants $f_{i}$ and $g_{i}$ (and similarly for the other operators). Computing the commutators $\left[\hat{a}_{\lambda}(\mathbf{k}^{\prime}),\hat{p}_{i}(\mathbf{k},\lambda)\right]$ and $\left[\hat{a}_{\lambda}(\mathbf{k}^{\prime}),\hat{p}^{\dagger}_{i}(-\mathbf{k},\lambda)\right]$ allows us to identify the constants,
@@ Appendix C Details of the diagonalisation
$\hat{b}_{\lambda}(\mathbf{k})$ $=\sum_{i}\beta_{i}^{*}\hat{p}_{i}(\mathbf{k},\lambda)-\widetilde{\beta}_{i}\hat{p}^{\dagger}_{i}(-\mathbf{k},\lambda),$ (44)
@@
Furthermore, inserting the expressions in Eqns. (44) into the commutators $\left[\hat{a}_{\lambda}(\mathbf{k}),\hat{a}^{\dagger}_{\lambda^{\prime}}(\mathbf{k}^{\prime})\right]=\delta_{\lambda\lambda^{\prime}}\delta(\mathbf{k}-\mathbf{k}^{\prime})$, etc., produces the relations
@@
$\sum_{i}|\beta_{i}|^{2}-|\widetilde{\beta}_{i}|^{2}$ $=1,$ (45)
@@
$(\omega_{i}-k)\alpha_{i}$ $=\kappa\left[-i(\beta_{i}-\widetilde{\beta}_{i})-\lambda(\gamma_{i}-\widetilde{\gamma}_{i})\right],$
$(\omega_{i}+k)\widetilde{\alpha}_{i}$ $=\kappa\left[i(\beta_{i}-\widetilde{\beta}_{i})-\lambda(\gamma_{i}-\widetilde{\gamma}_{i})\right],$
$\beta_{i}$ $=i\kappa\left(\frac{\alpha_{i}+\widetilde{\alpha}_{i}}{\omega_{i}-\widetilde{\omega}_{0}}\right),$
$\widetilde{\beta}_{i}$ $=i\kappa\left(\frac{\alpha_{i}+\widetilde{\alpha}_{i}}{\omega_{i}+\widetilde{\omega}_{0}}\right),$ (46)
$\gamma_{i}$ $=-\lambda\kappa\left(\frac{\alpha_{i}-\widetilde{\alpha}_{i}}{\omega_{i}-\widetilde{\omega}_{0}}\right),$
$\widetilde{\gamma}_{i}$ $=-\lambda\kappa\left(\frac{\alpha_{i}-\widetilde{\alpha}_{i}}{\omega_{i}+\widetilde{\omega}_{0}}\right).$
@@
$|\alpha_{i}|^{2}-|\widetilde{\alpha}_{i}|^{2}+|\beta_{i}|^{2}-|\widetilde{\beta}_{i}|^{2}+|\gamma_{i}|^{2}-|\widetilde{\gamma}_{i}|^{2}=1.$ (47)
@@
$n(\omega_{i})=\frac{k}{\omega_{i}}=\left(1-\frac{\alpha^{2}}{\omega_{i}^{2}-\omega_{0}^{2}}\right),$ (48)
@@
$n(\omega_{i})=\frac{k}{\omega_{i}}=-\left(1-\frac{\alpha^{2}}{\omega_{i}^{2}-\omega_{0}^{2}}\right),$ (49)
@@
where we have chosen the modes $\omega_{i}=\{\omega_{u},\omega_{p},\omega_{d}\}$ in descending order of oscillation frequency. Manipulating the algebratic relations in Eqns. (C) yields
@@
$\left|\alpha_{u/d}\right|^{2}$ $=\left(1+2\kappa^{2}\left[\frac{1}{(\omega_{u/d}-\widetilde{\omega}_{0})^{2}}-\frac{1}{(\omega_{u/d}+\widetilde{\omega}_{0})^{2}}\right]\right)^{-1}$ (50)
$\left|\widetilde{\alpha}_{p}\right|^{2}$ $=\left(-1+2\kappa^{2}\left[\frac{1}{(\omega_{p}-\widetilde{\omega}_{0})^{2}}-\frac{1}{(\omega_{p}+\widetilde{\omega}_{0})^{2}}\right]\right)^{-1}$ (51)
using Eqns. (C) and the normalisation condition Eq. (47). Importantly, we can note that this satisfies
@@
which must be the case in order to preserve the canonical commutation relations. The phases of $\alpha_{u/d}$ and $\widetilde{\alpha}_{p}$ can be chosen freely, and the remaining Hopfield coefficients follow from Eqns. (C).
Appendix D Diagonal form of the helicity
In this section, we provide details of how to obtain Eq. (29) in the main text, using the polariton coefficients obtained in the previous section. Let us express the total helicity, Eq. (9), in terms of the photon and matter annihilation and creation operators, yielding:
$\int{d^{3}r:\hat{h}:}$ $=\int d^{3}r\;\left[\frac{1}{2}\left(\mathbf{\widehat{A}}\cdot\mathbf{\widehat{B}}-\mathbf{\widehat{C}}\cdot\mathbf{\widehat{D}}\right)+\boldsymbol{\widehat{\Pi}}^{P}\cdot\mathbf{\widehat{M}}-\boldsymbol{\widehat{\Pi}}^{M}\cdot\mathbf{\widehat{P}}\right$]